OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6

OpenStudy (mathmale):

Ed: This is a great opportunity to apply the "point-slope formula for the equation of a straight line":\[y-y _{?}=m(x-x _{0)}\]

OpenStudy (mathmale):

as you are given a point on the line whose equation you want to find. But where does the necessary slope, m, come from? You're given the line y = -x + 6. But the slope of the desired line is not -1. What is it, and how do you obtain that slope?

OpenStudy (anonymous):

quick question what do i do with the -x

OpenStudy (anonymous):

choose or find? for perpendicular gradient, always use this formula:   y=mx+c y=1x+c substitute in the point as it is shared with both lines -2=1(-4)+c c=? then find your equation